When flipping a coin twice what is the probability for getting two heads given that there is at least one head already?

When flipping a coin twice what is the probability for getting two heads given that there is at least one head already?

The first is the Product Rule. This states that the probability of the occurrence of two independent events is the product of their individual probabilities. The probability of getting two heads on two coin tosses is 0.5 x 0.5 or 0.25.

What is the probability if you flip a coin twice and it lands on tails in both trials?

The odds of getting tails twice in a row are 1/2 * 1/2 = 1/4. So 25% of the time you’ll get heads twice in a row. But, if you’ve already flipped a coin 99 times, and gotten heads each time, then the odds of your next flip being heads is still just 50:50.

What is the probability that a coin will land on heads when flipped?

Suppose you have a fair coin: this means it has a 50% chance of landing heads up and a 50% chance of landing tails up. Suppose you flip it three times and these flips are independent. What is the probability that it lands heads up, then tails up, then heads up? So the answer is 1/8, or 12.5%.

What is the probability of flipping a coin 3 times and getting heads all 3 times?

0.125
Answer: If you flip a coin 3 times the probability of getting 3 heads is 0.125.

What is the probability of flipping heads 20 times in a row?

12
The probability of flipping a head after having already flipped 20 heads in a row is 12. Assuming a fair coin: The probability of 20 heads, then 1 tail is 0.520 × 0.5 = 0.5. The probability of 20 heads, then 1 head is 0.520 × 0.5 = 0.5.

What is the probability that when a fair coin is flipped six times in a row it lands heads up every time?

A coin has two possible outcomes: heads and tails. Thus there are 64 possible outcomes of the six flips and 1 of the 64 outcomes results in heads every time.

What is the probability of flipping 2 heads in a row?

There is a 1/4 chance of getting two heads in a row when tossing a coin twice.

What is the probability of a two headed coin?

The probabilities are now 2/3 and 1/3 for where the 2 headed coin is, after seeing one head. One head makes it twice as likely the coin is 2 headed, not a fair coin. I think the discussion is far too complicated even though the correct answer of 2 out of 3 chances has been given.

What is the probability of two fair coins being tossed?

The first coin must show Heads AND the second coin must show Heads. P (H,H) = P (H) AND P (H) ← ‘and’ means multiply P (H) × P (H) = 1 2 × 1 2 = 1 4

What makes a coin a double headed coin?

Therefore for the coin to be the double headed coin, the other side has to be a head of which Heads is 2 of the 3 equal possibilities. Those who give the answer of 50/50 are confusing possibility with chance. As there are 2 heads and only one tail as possibilities, the odds or chances are 2 out of 3 being heads.

What is the probability of obtaining a head?

Let H ≡ obtaining heads, A ≡ picking a two-headed coin, B ≡ picking a two-tailed coin, and C ≡ picking a fair coin. Then observe that the probability of obtaining a head is: The possible outcomes are: All these outcomes are equally likely. How many of these are heads and of those how many use the double headed coin? The answer is 2/3.